OpenStudy (anonymous):

What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374? 11 11.5 12 12.5

OpenStudy (anonymous):

do you have any formulas? look them up in wiki

OpenStudy (anonymous):

http://www.mathwords.com/a/arithmetic_sequence.htm

OpenStudy (anonymous):

The first term is a1, the common difference is d, and the number of terms is n. an = a1 + (n – 1)d plug in 9374= -13+ (43 – 1)d solve for d

OpenStudy (anonymous):

223.5? whatttt

OpenStudy (anonymous):

http://www.mathwords.com/a/arithmetic_series.htm here another formula

OpenStudy (anonymous):

use the series not sequence lol

OpenStudy (campbell_st):

well i think you have made a error.... the sum of an arithmetic series is \[S_{n} = \frac{n}{2}[2a + (n -1)d]\] use the information to solve for d \[9374 = \frac{43}{2}[2 \times -13 + (43 -1)d]\] the formula shown by @timo86m is for calculating a term in an arithmetic sequence

OpenStudy (anonymous):

9374 = 43/2*(2*(-13)+(43-1)*d) d=11

OpenStudy (anonymous):

Thank you both., I had the equation going ahhaa

OpenStudy (anonymous):

Might I ask one more?

OpenStudy (anonymous):

well i am going to bed :( sorry

OpenStudy (anonymous):

Oh okay.